Solve the following IVPs using Laplace Transform: 2) y + 4y + 3y = 3 ezt; y(0) = y(0) = 0
Answer #1


S-2 3 S-2 S-2 3 S-2 y+4y+3y = 3e24 Applying laplace transform on both sides L(y+4y+3y) = 2(322) 3 L(y)+44(y)+32(y) = (s